OpenStudy (peachy1):

Math 5 questions please help i am not sure about thease

rishavraj (rishavraj):

area of circle = pi*r^2 r = 4 so area = \[\pi \times 4^2\]

rishavraj (rishavraj):

@Peachy1 ????

rishavraj (rishavraj):

so u got the answer

rishavraj (rishavraj):

yup

rishavraj (rishavraj):

whts the radius ??? the diameter given is 18 ...so whts the radius ?? @Peachy1

rishavraj (rishavraj):

how/?

rishavraj (rishavraj):

18 is the diameter .wht is the radius ??

rishavraj (rishavraj):

@Peachy1 http://www.mathopenref.com/radius.html

rishavraj (rishavraj):

first of al,l study about radius and diameter

rishavraj (rishavraj):

u wil be able to solve by urself

rishavraj (rishavraj):

good :))) @Peachy1